A General Power bond carries a coupon rate of 9.1%, has 9 years until maturity, and sells at a yield to maturity of 8.1%. (Assume annual interest payments.) a. What interest payments do bondholders receive each year? Interest payments S 91 b. At what price does the bond sell? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Price c. what will happen to the bond price if the yield to maturity falls to 71%? (Do not round intermediate calculations Round your answer to 2 decimal places Price wil by
Answer #1
a) Interest payments received each year = $1000 * 9.1% = $91
b) Price of the bond: coupon * PVIFA(yield , n) + par value * PVIF(yield , n)
Price of the bond= 91 * PVIFA(8.1% , 9) + 1000 * PVIF(8.1% , 9)
Price of the bond= 91 * 6.2210 + 1000 * 0.4961 = $1062.21
c) Price of the bond: coupon * PVIFA(yield , n) + par value * PVIF(yield , n)
Price of the bond= 91 * PVIFA(7.1% , 9) + 1000 * PVIF(7.1% , 9)
Price of the bond= 91 * 6.4876 + 1000 * 0.5394 = $1129.77
The price will increase by $67.56 (ie. 1129.77 - 1062.21)